OpenStudy (anonymous):

Dividing Rationale Expressions Simplify: 4x^2/16x^3 (four x square divided by sixteen x cube)

OpenStudy (anonymous):

\[\frac{4x^2}{16x^3}=\frac{4 \times x \times x}{4 \times 4 \times x \times x \times x}\]

OpenStudy (anonymous):

You have common factors of 4 and (x times x) above and below. You can cancel these out to get\[\frac{1}{4x}\]

OpenStudy (anonymous):

16x^2y/24xy^3

OpenStudy (anonymous):

2x/3y^2 is this correct?

OpenStudy (anonymous):

\[\frac{16x^2y}{24xy^3}=\frac{2x}{3y^2}\]Yes, that's correct.